A volume of 0.50 L of water at 16°C is put into an aluminum ice cube tray of mass 0.250 kg at the same temperature. How much energy must be removed from this system by the refrigerator to turn the water into ice at -8.0°C?


1
Expert's answer
2021-02-16T10:30:32-0500

Solution:


Specific heat capacity of water is 4179.6 ["\frac{J}{(kg\cdot K)}" ]

The enthalpy of fusion is 333.55 ["\frac{J}{g}"]

Specific heat of ice is 2.04 ["\frac{J}{(g\cdot K)}"]

0.5 liters = 0.5 kg

0.5 kg * 4179.6 "\frac{J}{(kg\cdot K)}" * (16 - 0) K + 0.5 kg * 333.55 "\frac{J}{g}"   +  0.5 kg * 2.04 "\frac{J}{(g\cdot K)}" * (0 - (-8)) K =>

0.5 * 4179.6 * 16 J + 0.5 * 333.55 kJ + 0.5 * 2.04 * 8 kJ =>

0.5 * (4.1796 * 16 + 333.55 + 2.04 * 8) kJ =>

0.5 * 416.7436 kJ => 208.3718 kJ


Specific heat of Aluminum is 0.9 "(\frac{J}{g\cdot K})"


E= 0.500*(4.186*16+333+2.09*8) = 208.3 kjoule

Ea = 0.25*0.9*(16+8) = 5.3 kjoule

E = Ew+E= 208.3 + 5.3 = 213.6 kjoule =214 kJ



Answer:


214 kJ